OpenStudy (anonymous):

Nina's monthly budget is $2,250. Every month she makes the following payments: $175 for insurance $129 for utilities $283 for taxes Which is the highest house payment she can afford? A. House A - monthly payment: $1,580.89 B. House B - monthly payment: $1,790.25 C. House C - monthly payment: $1,989.23

OpenStudy (anonymous):

2250-(all the utilities)=2250-(283+129+175)=2250-(587)

OpenStudy (anonymous):

can you calculate the rest ;)

OpenStudy (anonymous):

So My Answer Is A Right

OpenStudy (anonymous):

yes i guess

OpenStudy (anonymous):

What Youu Mean I Guess

OpenStudy (anonymous):

i mean yes it must be

OpenStudy (anonymous):

Thannks ")

OpenStudy (anonymous):

:) YW and you ain't bad you are beautiful :)